# One-Step Offline Distillation of Diffusion-based Models via Koopman Modeling

### v2

@@ 5 Theoretical Properties of KDM
The proof builds on the analytic formulation of $\Phi$, allowing a convergent multivariate Taylor expansion [71], and on exact finite-dimensional Koopman operators for polynomial dynamics [26], see App. A.
@@
Let $\Phi_{t}:\mathbb{R}^{n}\to\mathbb{R}^{n}$ be the reverse diffusion flow of a trained model (from time $T$ to 00), and let $\mathcal{O}\subset L^{2}(\mathbb{R}^{n})$ be a finite-dimensional subspace such that:
@@
The proof builds on the observation that diffusion model dynamics tend to preserve semantic similarity between inputs and outputs. By leveraging the Koopman operator’s spectral properties, we construct a finite-dimensional space of functions that remains stable under the model’s evolution. Within this space, the Koopman coordinates $\xi$ provide a locally bi-Lipschitz representation, while the reverse diffusion map $\Phi_{T}$ is Lipschitz continuous. Together, these properties ensure that semantically similar inputs remain close throughout the diffusion process. Note that the theorem holds also for $\xi:=\text{id}$, i.e., the identity map [48, 15]. The full proof is given in App. B.
@@ 6 Experiments
We begin by broadening the investigation of the emergent structure in the dynamics that map noise to images, and its presence in KDM (Sec. 6.1). We then extensively evaluate our method on the offline distillation benchmark (Sec. 6.2), including analyses of computational complexity, data and parameter scaling, and denoiser-agnostic applicability. Additionally, an ablation study of the loss functions is provided in App. D.1. We also provide a code repository of KDM in the supplementary materials.
@@
We follow the protocol of GET [19] for generating noisy-clean pairs and for evaluation, using the CIFAR-10 [37], FFHQ 64$\times$64 [30], and AFHQv2 64$\times$64 [11] datasets. Our training setup matches that of PD [58] and GET [19]. For evaluation, we report image quality using Fréchet Inception Distance (FID) and Inception Score (IS), computed over 50k samples. Both encoders $E_{\varphi}$ and $E_{\phi}$, as well as the decoder $D_{\psi}$, are implemented as compact UNets with reduced output channels to ensure similar parameters size to the other single-step methods. The discriminator consists of four simple Conv2D layers. Full implementation details are provided in App. E.3.
@@ 6.1 Observations of Emergent Latent Structure
Figure 3: Visualization of vicinity comparison across different noise levels $\sigma$. Left: EDM [28]. Right: our method. Each image corresponds to a small perturbation around the original noise sample.
@@
Extending the 2D toy experiment from Sec. 4, we now explore the high-dimensional FFHQ 64$\times$64 dataset. Since dimensionality reduction can obscure fine-grained patterns, we instead analyze the local neighborhood of a single latent sample $x_{T}\sim\mathcal{N}(0,I_{n})$. We generate perturbed versions $\hat{x}_{T}=x_{T}+\varepsilon\cdot\sigma$ with $\varepsilon\sim\mathcal{N}(0,I_{n})$, and normalize each $\hat{x}_{T}$ using its own mean and standard deviation. This setup tests whether semantic structure emerges locally around $x_{T}$. We vary $\sigma\in\{0,0.05,\dots,0.60\}$ and visualize the outputs in Fig. 3 (left). For all $\sigma<0.3$, the images remain semantically aligned with the original image (when $\sigma=0$), indicating local smoothness and manifold adherence. As $\sigma$ grows, semantics diverge but retain global coherence even at $\sigma=0.60$. The same experiment on KDM (Fig. 3, right) reveals similar behavior, affirming that KDM preserves the local semantic structure of the original EDM. Our results show that both models organize latent space into meaningful neighborhoods, even in high-dimensional settings.
@@
Beyond preserving local structure, KDM also approximately reproduces the noise-to-image mapping of EDM. When starting from the same noise vector $x_{T}\sim\mathcal{N}(0,I_{n})$, sampled using seeds not seen during training, both EDM (with 79 NFEs) and KDM (with a single NFE) generate perceptually similar outputs (see Fig. 4). This suggests that KDM generalizes well beyond its training data and effectively distills the mapping from noise to data. We observe this consistency across multiple datasets and include further qualitative results in App. D.2. To quantify this, we generate 50k samples using both methods and compare outputs using LPIPS and MSE. KDM achieves scores of 0.13 (LPIPS) and 0.009 (MSE), in contrast to 0.48 and 0.57 when comparing randomly permuted images, demonstrating the similarity between the models’ generation. To complement quantitative metrics, we conduct a human study where participants choose the more realistic image collage between our model and EDM (Fig. 4; details in App. E.4). Evaluators showed no clear preference: $55\%$ of 300 choices favored our model, highlighting its strong fidelity.
Figure 4: Sampled images using 79 NFEs with EDM [28] (left), and using 1 NFE with KDM (right).
@@
We analyze a recurring failure mode in both the teacher and distilled models: the generation of outlier samples that fall outside the designated checkerboard regions. These out-of-distribution points often appear in the final outputs and reflect a limitation in capturing the intended data distribution. To investigate their origin, we trace these outliers back through the generation process and identify two primary sources: (1) the tails of the Gaussian prior and (2) regions near decision boundaries between adjacent checkerboard cells. Fig. 5(c) shows this behavior for the distilled model (black crosses denote outliers); additional comparisons with the teacher are provided in App. E.6. These findings suggest that the model’s dynamics are more error-prone in high-uncertainty regions, offering potential directions for improvement. Integrating more robust sampling techniques or leveraging self-supervised signals may help mitigate such failures. This insight could also inform future efforts in uncertainty estimation.
@@ 6.2 One-step Generation Results
We evaluate our method on the standard offline distillation benchmark introduced in [19], which tests the ability to distill a pre-trained EDM on the CIFAR-10 dataset. While EDM generates samples in 35 NFEs, our approach achieves the results in a single NFE. Tab. 3 reports unconditional generation results, where our method achieves state-of-the-art (SOTA) performance, surpassing the previous offline baseline in both FID and IS. In particular, we observe $\approx 25\%$ reduction in FID over the best prior method. Tab. 3 shows results for class-conditional generation, where our model yields $\approx 40\%$ improvement in FID and superior IS. Although offline distillation offers advantages in efficiency, privacy, and flexible deployment [17, 19], a performance gap with online methods remains. This may partly be because online approaches leverage the full generation trajectory, whereas, in the offline pairwise setup, methods only observe two endpoints, placing them at a natural disadvantage. Our results demonstrate a substantial step toward closing this gap. Furthermore, unlike prior methods that rely on architectural assumptions [20, 39], our approach is denoiser-agnostic and successfully distills both EDM and flow matching (FM) models (see Tab. 3), demonstrating broad applicability. Finally, we implemented a factorized KDM (KDM-F). Since every matrix $C_{\eta}$ is diagonalizable up to an arbitrarily small perturbation of the entries [2], we can write: $C_{\eta}=P\Lambda P^{-1}$ where $P\in\mathbb{C}^{d\times d}$ is an invertible matrix and $\Lambda=\text{diag}(\lambda_{1},\dots,\lambda_{d})\in\mathbb{C}^{d\times d}$. While enabling efficient implementation of spectral penalties on $\Lambda$ (see complexity analysis below), this approach preserves strong performance and accuracy.
@@
Models (($\downarrow$( ↓) / Metrics ($\rightarrow$) NFE $\downarrow$ FID $\downarrow$ IS $\uparrow$
@@
DDPM [24] 1000 3.17 9.46
Score SDE [63] 2000 2.2 9.89
EDM [30] 35 2.04 9.84
@@
Flow Matching (Diffusion) [40] 183 8.06 -
Flow Matching (OT) [70] 50 4.94 -
2-rf++ [39] 1 3.07 -
@@
KD [44] 1 9.36 8.36
PD [58] 1 9.12 -
DFNO [79] 1 4.12 -
iCD-EDM (LPIPS) [64] 1 2.83 9.54
CTM (LPIPS) [64] 1 1.98 -
ECM [20] 1 4.54 -
@@
Models (($\downarrow$( ↓) / Metrics ($\rightarrow$) NFE $\downarrow$ FID $\downarrow$ IS $\uparrow$
@@
BigGAN [7] 1 14.73 9.22
StyleGAN2-ADA [29] 1 2.42 10.14
@@
Guided Dist. ($w=0.3$[49] 1 7.34 8.90
ECM [20] 1 3.81 -
@@
Table 3: Generative performance on unconditional generation on FFHQ 64$\times$64 and AFHQv2 64$\times$64. *Indicates our reproduction.
Models (($\downarrow$( ↓) / Metrics ($\rightarrow$) NFE $\downarrow$ FID $\downarrow$ IS $\uparrow$
FFHQ 64$\times$64
@@
To assess the scalability of KDM to higher-resolution data, we evaluate our method on FFHQ and AFHQv2, distilling each into a single-step generator and assessing performance both quantitatively (Tab. LABEL:table:additional_uncond_datasets) and qualitatively (App. D.2). Our method yields FID scores approximately 2.5$\times$ higher than EDM, similar to the degradation observed on CIFAR-10. Due to the high computational cost, we omit GET results: training a single model takes 5–7 weeks on an A6000 GPU, which is technically infeasible in our computational resources. We also exclude ImageNet, as large-scale training with EDM [28] is computationally infeasible for us.
@@
Offline distillation inherently supports architectural independence from the teacher model, allowing flexible scaling of the student to meet diverse deployment needs [19, 20]. To assess this, we evaluate our method against GET across three model sizes—20M, 38M, and 84M parameters (Fig. 5A). Our method consistently outperforms GET in FID scores at all scales, with especially notable gains at higher capacities (e.g., 4.89 vs. 7.19 at 84M), demonstrating its effectiveness for both lightweight and high-capacity settings. In parallel, we investigate how performance scales with training data size, a key factor in offline distillation. Following GET [19], we train on 250K, 500K, and 1M samples and report FID scores on unconditional CIFAR-10 in Fig. 5B. Results show that more data yields better generation quality. For instance, increasing the dataset from 250K to 1M samples results in a $\approx 33\%$ FID reduction, showing the importance of data scale. While our method remains data-efficient, additional data can enhance performance.
@@
Complexity Analysis.
Offline distillation can offer advantages over online methods, both memory-wise and time-wise [19]. KDM further improves over GET. As shown in Tab. 4, it matches GET in parameter count but achieves a 4×4$\times$4 × speedup per training iteration and over 8×8$\times$8 × faster sampling. Notably, KDM-F (App. C) maintains this efficiency even with spectral regularization. These results underscore the practical scalability of our approach.
@@ 7 Discussion
In this work, we present two key observations. First, we reinterpret the distillation process through the lens of dynamical systems theory. Second, we uncover the semantical structure in the mapping from latent space to data space in trained diffusion dynamics. Motivated by these insights, we propose a Koopman-based framework for distilling diffusion models, bridging operator theory and generative modeling. We theoretically establish the existence of a Koopman operator that captures the underlying generative process, and demonstrate, both theoretically and empirically, that our method preserves its structure. This perspective also offers insights into potential failure modes. Beyond its conceptual contributions, our approach is practical and efficient, achieving state-of-the-art single-step offline distillation. We further analyze its robustness and scalability. We hope this work lays the foundation for future directions, including leveraging richer training signals and extending distillation techniques to advanced text-to-image generative models. See App. F for further discussion.

### v3

@@ 5 Theoretical Properties of KDM
The proof builds on the analytic formulation of $\Phi$, allowing a convergent multivariate Taylor expansion trefethen2019approximation , and on exact finite-dimensional Koopman operators for polynomial dynamics iacob2023finite , see App. A.
@@
Let $\Phi_{t}:\mathbb{R}^{n}\to\mathbb{R}^{n}$ be the reverse diffusion flow of a trained model (from time $T$ to 0), and let $\mathcal{O}\subset L^{2}(\mathbb{R}^{n})$ be a finite-dimensional subspace such that:
@@
The proof builds on the observation that diffusion model dynamics tend to preserve semantic similarity between inputs and outputs. By leveraging the Koopman operator’s spectral properties, we construct a finite-dimensional space of functions that remains stable under the model’s evolution. Within this space, the Koopman coordinates $\xi$ provide a locally bi-Lipschitz representation, while the reverse diffusion map $\Phi_{T}$ is Lipschitz continuous. Together, these properties ensure that semantically similar inputs remain close throughout the diffusion process. Note that the theorem holds also for $\xi:=\text{id}$, i.e., the identity map mauroy2020introduction ; dogra2020optimizing . The full proof is given in App. B.
Importantly, we observe in practice an average spectral norm of the Koopman operator $C_{T}$ of $0.88\pm 0.18$, indicating a reasonably tight bound and supporting the practical utility of our theory. Moreover, our framework enables explicit control over this norm via eigenvalue penalties on the Koopman operator. As detailed in App. C, we factorize $C_{T}$ to permit direct constraints on its spectrum (e.g., radius or banding), providing a principled regularizer on $\|C_{T}\|$.
@@ 6 Experiments
We begin by further investigating the emergent dynamics structure that map noise to images, and its presence in KDM (Sec. 6.1). We then extensively assess our method on the offline distillation benchmark (Sec. 6.2), including analyses of computational complexity, data and parameter scaling, and denoiser-agnostic utility. Additionally, an ablation study of the loss functions and adversarial losses stability is provided in App. D.1. Our code is in https://github.com/azencot-group/KDM.
@@
We follow the protocol of GET geng2023one for generating noisy-clean pairs and for evaluation, using the CIFAR-10 krizhevsky2009learning , FFHQ $\times64$ karras2019style , and AFHQv2 $\times64$ choi2020stargan datasets. Our training setup matches that of PD salimans2022progressive and GET geng2023one . For evaluation, we report image quality using Fréchet Inception Distance (FID) and Inception Score (IS), computed over 50k samples. Both encoders $E_{\varphi}$ and $E_{\phi}$, as well as the decoder $D_{\psi}$, are implemented as compact UNets with reduced output channels to ensure a similar parameter size to the other single-step methods. The discriminator consists of four simple Conv2D layers. Full implementation details are provided in App. E.3.
@@ 6.1 Observations of Emergent Latent Structure
Figure 3: Visualization of vicinity comparison across different noise levels $\sigma$. Left: EDM karras2022elucidating . Right: our method. Each image corresponds to a small perturbation around the original noise sample.
@@
Extending the 2D toy experiment from Sec. 4, we now explore the high-dimensional FFHQ $\times64$ dataset. Since dimensionality reduction can obscure fine-grained patterns, we instead analyze the local neighborhood of a single latent sample $x_{T}\sim\mathcal{N}(0,I_{n})$. We generate perturbed versions $\hat{x}_{T}=x_{T}+\varepsilon\cdot\sigma$ with $\varepsilon\sim\mathcal{N}(0,I_{n})$, and normalize each $\hat{x}_{T}$ using its own mean and standard deviation. This setup tests whether semantic structure emerges locally around $x_{T}$. We vary $\sigma\in\{0,0.05,\dots,0.60\}$ and visualize the outputs in Fig. 3 (left). For all $\sigma<0.3$, the images remain semantically aligned with the original image (when $\sigma=0$), indicating local smoothness and manifold adherence. As $\sigma$ grows, semantics diverge but retain global coherence even at $\sigma=0.60$. The same experiment on KDM (Fig. 3, right) reveals similar behavior, affirming that KDM preserves the local semantic structure of the original EDM. Our results show that both models organize latent space into meaningful neighborhoods, even in high-dimensional settings.
@@
Beyond preserving local structure, KDM also approximately reproduces the noise-to-image mapping of EDM. When starting from the same noise vector $x_{T}\sim\mathcal{N}(0,I_{n})$, sampled using seeds not seen during training, both EDM (with 79 NFEs) and KDM (with a single NFE) generate perceptually similar outputs (see Fig. 4). This suggests that KDM generalizes well beyond its training data and effectively distills the mapping from noise to data. We observe this consistency across multiple datasets and include further qualitative results in App. D.3. To quantify this, we generate 50k samples using both methods and compare outputs using LPIPS and MSE. KDM achieves scores of 0.13 (LPIPS) and 0.009 (MSE), in contrast to 0.48 and 0.57 when comparing randomly permuted images, demonstrating the similarity between the models’ generation. To complement quantitative metrics, we conduct a human study where participants choose the more realistic image collage between our model and EDM (Fig. 4; details in App. E.4). Evaluators showed no clear preference: $55\%$ of 300 choices favored our model, highlighting its strong fidelity.
Figure 4: Sampled images using 79 NFEs with EDM karras2022elucidating (left), and using 1 NFE with KDM (right).
@@
We analyze a recurring failure mode in both the teacher and distilled models: the generation of outlier samples that fall outside the designated checkerboard regions. These out-of-distribution points often appear in the final outputs and reflect a limitation in capturing the intended data distribution. To investigate their origin, we trace these outliers back through the generation process and identify two primary sources: (1) the tails of the Gaussian prior and (2) regions near decision boundaries between adjacent checkerboard cells. Fig. 5C shows this behavior for the distilled model (black crosses denote outliers); additional comparisons with the teacher are provided in App. E.7. These findings suggest that the model’s dynamics are more error-prone in high-uncertainty regions, offering potential directions for improvement. Integrating more robust sampling techniques or leveraging self-supervised signals may help mitigate such failures. This insight could also inform future efforts in uncertainty estimation.
@@ 6.2 One-step Generation Results
We evaluate on the standard offline distillation benchmark of (geng2023one, ), which distills a pre-trained EDM on CIFAR-10. Whereas the EDM teacher samples with 35 function evaluations (NFEs), our student attains comparable quality with a single NFE. We also adapt two recent SOTA baselines, IMM (zhou2025inductive, ) and RF-2++ (lee2024improving, ), to the same strictly offline setting, where only teacher-generated data are available and the original model weights (and online queries) are not accessible (the RF-2++ variant corresponds to its “w/o pre-train” configuration). All methods are evaluated under identical protocols. Tab. 3 reports unconditional generation results, where our method achieves highly competitive performance, surpassing the previous offline baseline in both FID and IS. In particular, we observe $\approx 25\%$ reduction in FID over GET. Tab. LABEL:table:conditional-cifar shows results for class-conditional generation, where our model yields $\approx 40\%$ improvement in FID and superior IS compared to GET. Although offline distillation offers advantages in efficiency, privacy, and flexible deployment fernandez2023privacy ; geng2023one , a performance gap with online methods remains. This may partly be because online approaches leverage the full generation trajectory, whereas, in the offline pairwise setup, methods only observe two endpoints or just the generated data, placing them at a natural disadvantage. Our results demonstrate a substantial step toward closing this gap.
Furthermore, unlike prior methods that rely on architectural assumptions geng2024consistency ; lee2024improving , our approach is denoiser-agnostic and successfully distills both EDM and flow matching (FM) models (see Tab. 3), demonstrating broad applicability. Finally, we implemented a factorized KDM (KDM-F). Since every matrix $C_{\eta}$ is diagonalizable up to an arbitrarily small perturbation of the entries axler2015linear , we can write: $C_{\eta}=P\Lambda P^{-1}$ where $P\in\mathbb{C}^{d\times d}$ is an invertible matrix and $\Lambda=\text{diag}(\lambda_{1},\dots,\lambda_{d})\in\mathbb{C}^{d\times d}$. While enabling efficient implementation of spectral penalties on $\Lambda$ (see complexity analysis below), this approach preserves strong performance and accuracy.
@@
Models $(\downarrow$) / Metrics ($\rightarrow$) NFE $\downarrow$ FID $\downarrow$ IS $\uparrow$
@@
DDPM ho2020denoising 1000 3.17 9.46
Score SDE song2020denoising 2000 2.2 9.89
EDM karras2019style 35 2.04 9.84
@@
Flow Matching (Diffusion) lipman2022flow 183 8.06 -
Flow Matching (OT) tong2023improving 50 4.94 -
2-rf++ lee2024improving 1 3.07 -
@@
PD salimans2022progressive 1 9.12 -
DFNO zheng2023fast 1 4.12 -
iCD-EDM (LPIPS) song2023improved 1 2.83 9.54
CTM (LPIPS) song2023improved 1 1.98 -
ECM geng2024consistency 1 4.54 -
@@
Models $(\downarrow$) / Metrics ($\rightarrow$) NFE $\downarrow$ FID $\downarrow$ IS $\uparrow$
@@
BigGAN brock2018large 1 14.73 9.22
StyleGAN2-ADA karras2020training 1 2.42 10.14
@@
Guided Dist. (w=0.3w=0.3meng2023distillation 1 7.34 8.90
ECM geng2024consistency 1 3.81 -
@@
Table 3: Generative performance on unconditional generation on FFHQ $\times64$ and AFHQv2 $\times64$. *Indicates our reproduction.
Models $(\downarrow$) / Metrics ($\rightarrow$) NFE $\downarrow$ FID $\downarrow$ IS $\uparrow$
FFHQ $\times64$
@@
[…]
@@
Offline distillation inherently supports architectural independence from the teacher model, allowing flexible scaling of the student to meet diverse deployment needs geng2023one ; geng2024consistency . To assess this, we evaluate our method against GET across three model sizes—20M, 38M, and 84M parameters (Fig. 5A). Our method consistently outperforms GET in FID scores at all scales, with especially notable gains at higher capacities (e.g., 4.89 vs. 7.19 at 84M), demonstrating its effectiveness for both lightweight and high-capacity settings. In parallel, we investigate how performance scales with training data size, a key factor in offline distillation. Following GET geng2023one , we train on 250K, 500K, and 1M samples and report FID scores on unconditional CIFAR-10 in Fig. 5B. Results show that more data yields better generation quality. For instance, increasing the dataset from 250K to 1M samples results in a $\approx 33\%$ FID reduction, showing the importance of data scale. While our method remains data-efficient, additional data can enhance performance.
Complexity analysis.
Offline distillation can offer advantages over online methods, both memory-wise and time-wise (see App. E.5). KDM further improves over GET. As shown in Tab. 4, it matches GET in parameter count but achieves a $4\times$ speedup per training iteration and over $8\times$ faster sampling. Notably, KDM-F (App. C) maintains this efficiency even with spectral regularization. Specifically, we evaluate KDM and KDM-F with an additional eigen-loss (EL) term, observing that KDM-F + EL requires training time comparable to KDM (without EL), whereas KDM + EL incurs a longer training duration. These results underscore the practical scalability of our approach.
@@ 7 Discussion
In this work, we present two key observations. First, we reinterpret the distillation process through the lens of dynamical systems theory. Second, we uncover the semantical structure in the mapping from latent space to data space in trained diffusion dynamics. Motivated by these insights, we propose a Koopman-based framework for distilling diffusion models, bridging operator theory and generative modeling. We theoretically establish the existence of a Koopman operator that captures the underlying generative process, and demonstrate, both theoretically and empirically, that our method preserves its structure. This perspective also offers insights into potential failure modes. Beyond its conceptual contributions, our approach is practical and efficient, achieving highly competitive single-step offline distillation. We further analyze its robustness and scalability. We hope this work lays the foundation for future directions, including leveraging richer training signals and extending distillation techniques to advanced text-to-image generative models. See App. F for further discussion.